# RSHazeDiff: A Unified Fourier-aware Diffusion Model for Remote Sensing Image Dehazing

### v1

@@ IV-C Global Compensated Learning
It is significant that the image restoration methods can handle unlimited-size images in practice. To achieve unlimited-size image restoration, we adopt the patch-based processing strategy that divides images into multiple fixed-size patches and then deals with each patch independently. However, this patch-based strategy may produce artifacts with the ignorance of global dependency information in all patches. To address this, we propose an effective GCL module that adopts FFB as the basic U-net architecture unit to capture all patches’ global dependency information. After learning global feature representations, the designed fusion module fuses the local patch-based image restoration process and GCL module in a pixel-channel attention collaborative manner. The detailed structure of GCL module and the fusion module are illustrated in Fig. 2.
@@ IV-D Phased Training Strategy
Apart from the above mentioned loss function that is optimized over local patch, we also exploit the perceptual loss to global compensated learning module for constraining the whole image. The perceptual loss function can measure high-level semantic/perceptual similarity between the output and ground truth by extracting the multi-layer features from pre-trained network VGG-19 [56]. By minimizing perceptual loss, the restored output $f_{2}$ has more similar texture and structure features with ground-truth images.
To be special, we hierarchically extract the latent feature maps from VGG-19 model to measure the image similarities between the output $f_{2}$ and ground-truth images $z$, which can be expressed as:
@@
where $F_{i}\left(\cdot\right)$ represents the $i$-th latent features obtained by the feature extraction network. $\omega_{i}$ denotes the weight coefficient of the $i$-th feature representation. Furthermore, we also employ pixel-level $\mathcal{L}_{1}$ loss to assess the diversity between $f_{2}$ and $z$. Hence, the loss function $\mathcal{L}_{global}$ in global compensated learning branch is shown as:
@@ V Experiments
In this section, the implementation details in our experiments are first described. Then, the comparisons with state-of-the-art approaches on the synthetic and real-world datasets are performed. Finally, we conduct ablation studies to demonstrate the effectiveness of core components in RSHazeDiff.
@@ V-A Implementation Details
Dataset. We evaluate the performance of RSHazeDiff on three RS hazy datasets, i.e., the light hazy image dataset (LHID), the dense hazy image dataset (DHID) [19] and expanded remote sensing image cloud removing dataset (ERICE) [19]. LHID and DHID are two synthetic datasets with light and dense haze, and each image has 512$\times$512 pixels. LHID contains 30,517 pairs of hazy RS images for training and 500 pairs for testing (i.e., LHID-A or LHID-B). In our experiments, we merge LHID-A and LHID-B, and test networks on LHID with 1,000 image pairs. Based on the atmospheric scattering model, the ambient light $A\in\left[0.7,1\right]$ and 500 different real-world transmission maps are randomly sampled to generate DHID dataset. DHID dataset contains 14,990 dense hazy images, with 14,490 for training and 500 for testing. The ERICE is a real-world dataset expanded on RICE1 [57], which are collected in some scene categories from Google Earth (i.e., mountain, river, desert, snowberg, forest, farmland, residential). In ERICE dataset, 1,850 and 150 images are regarded as training and test sets, each with a resolution of 256$\times$256 pixels.
TABLE I: Quantitative comparisons (Average PSNR/SSIM/CIEDE2000) with 11 state-of-the-art dehazing approaches on the synthetic LHID and DHID datasets. The $\uparrow$ means higher is better while $\downarrow$ means lower is better. The best results are written in boldface
@@
DCP [3] TPAMI’10 21.12 0.818 9.161 18.92 0.824 9.867
AOD-Net [12] ICCV’17 22.24 0.842 7.122 15.37 0.718 14.272
GridDehazeNet [14] ICCV’19 26.64 0.883 4.597 25.51 0.868 5.053
IDeRs [6] INS’19 14.68 0.627 19.563 13.33 0.575 19.942
FFA-Net [15] AAAI’20 27.79 0.887 4.424 25.05 0.880 5.590
PSD [16] CVPR’21 23.33 0.855 6.998 23.30 0.853 6.911
FCTF-Net [18] GRSL’21 25.85 0.876 5.423 17.21 0.739 12.794
DCIL [19] TGRS’22 28.12 0.898 5.180 26.95 0.896 4.877
Trinity-Net [35] TGRS’23 24.78 0.867 6.737 25.86 0.878 5.479
Focal-Net [17] ICCV’23 29.39 0.899 4.012 25.73 0.887 5.313
RSHazeDiff Ours 29.65 0.905 3.875 27.91 0.900 3.945
@@
Training Details. All experiments are implemented by PyTorch [58] framework with NVIDIA GeForce RTX 3090 GPU. In the training of Fourier-aware conditional diffusion model, we adopt the Adam optimizer and the initial learning rate of 0.00002. The batch and patch size are set as 4 and 64$\times$64. We adopt PTS to train the model until it converges. Concretely, we first train the model with noise estimation constraint for 1,500,000 iterations, and then refine it with reconstruction constraint for 500,000 iterations in LHID. Similarly, we first train the model for 1,000,000 iterations and then refine it for 500,000 iterations in DHID. To expedite the sampling process, we adopt the implicit sampling strategy [59] in which the total number of time steps and implicit sampling steps are set to 1,000 and 10. Furthermore, we use the exponential moving average (EMA) strategy [60] to eliminate slight color shift problem. During the training of the global compensated learning module, we use the Adam optimizer with the batch size of 8 and the initial learning rate of 0.001.
@@
Evaluation Settings. To demonstrate the superiority of RSHazeDiff, we qualitatively and quantitatively compare it against ten representative SOTA dehazing algorithms, including DCP [3], AOD-Net [12], GridDehazeNet [14], IDeRs [6], FFA [15], PSD [16], FCTF-Net [18], DCIL [19], Trinity-Net [35], and Focal-Net [17].
@@
The three well-known full-referenced quantitative evaluation metrics Peak Signal to Noise Ratio (PSNR), Structural Similarity Index (SSIM) [61] and CIEDE2000 [62] are adopted for dehazing performance evaluation. PSNR index assesses the error between the dehazed images and the corresponding ground-truth ones pixel by pixel, while SSIM measures image similarity regarding brightness, contrast, and structure. Larger values of PSNR and SSIM indicate better results. CIEDE2000 indicator is employed to evaluate the color distinction between the dehazed image and its haze-free counterpart. A smaller value of CIEDE2000 indicates better performance.
@@ V-B Comparison with State-of-the-art Methods
Results on Synthetic Dataset. Table I reports the quantitative performance of eleven dehazing algorithms on synthetic LHID and DHID test datasets. To make fair comparisons, we re-train all compared methods on synthetic datasets according to the implementation details of their papers. Note that we directly re-train the whole network of DCIL on DHID dataset without the phased learning strategy [19]. As observed, our RSHazeDiff can achieve the highest PSNR and SSIM values and the lowest CIEDE2000 values than other representative SOTA dehazing algorithms, which indicates that our RSHazeDiff is able to better restore the texture details and color information from hazy RS images.
@@
Fig. 5 presents the visual comparisons against several state-of-the-art dehazing approaches on DHID datasets. We can observe that AOD-Net tends to darken the whole images and produce noticeable Halo artifacts, degrading the overall visibility of the dehazed images. FFA-Net and Focal-Net cannot remove haze effectively when dealing with dense hazy images, leading to some haze residuals. In addition, the dehazing results of FFA-Net suffer from slight color distortion. Although DCIL can remove a majority of haze to a certain extent, it may distort some texture details of input images and fails to completely remove the haze in some regions. Trinity-Net cannot produce pleasing dehazing results due to haze residuals and serious color distortion. Compared with above SOTA algorithms, our RSHazeDiff can generate much clearer haze-free images with realistic color fidelity and finer texture details.
@@
Results on Real-World Hazy Images. To further validate the generalization ability on real-world images, we evaluate our method against other state-of-the-art dehazing approaches on ERICE dataset. The ERICE dataset contains some images captured in thick haze scenarios. Table II exhibits the quantitative comparison results over eleven representative dehazing algorithms. As reported, the proposed method surpasses other dehazing algorithms in terms of PSNR and SSIM, indicating that the images restored by RSHazeDiff are most similar to its corresponding ground-truth images. Although CIEDE2000 value of RSHazeDiff is not the best, our method wins the second place among eleven competitive dehazing algorithms. The quantitative results verify that the superiority of RSHazeDiff on real-world RS images.
@@
TABLE II: Average PSNR, SSIM and CIEDE2000 of various different approaches on the real-world ERICE dataset. red and blue colors are used to indicate the $1^{st}$ and $2^{nd}$ ranks, respectively
@@ V-D Ablation Study
Effect of different components in RSHazeDiff. Our method shows dehazing superiority both quantitatively and qualitatively over eleven representative SOTAs. Furthermore, we perform an ablation study to analyze the effectiveness of the proposed components, including the phased training strategy, Fourier-aware iterative refinement, and global compensated learning module.
@@
For fair comparisons, all variants are trained in the same way as before and tested on DHID dataset. The performances of these variants can be found in Table III and Fig. 1.
TABLE III: Ablation study of different components on DHID dataset.
@@
As shown in Table III and Fig. 1, each component of RSHazeDiff can boost the dehazing performance. When applying the phased training strategy, the model achieves advanced performance that has fewer haze residuals and higher metric values over baseline model. Furthermore, the introduction of FIR and GCL module also have brought performance improvement of model in three indicators. In short, the full model performs favorably against the other variants both quantitatively and qualitatively, demonstrating that all the components of RSHazeDiff are beneficial and necessary for effective haze removal.
Effect of feature fusion module in GCL. Since simple addition or concatenation operations cannot fuse two level features well, we employ two attention modules to enhance the feature fusion capability in a pixel-channel collaborative manner. To explore the performance of different fusion modes in the feature fusion module, we perform an ablation study by setting PA and CA in parallel or in serial, or swapping the addition and concatenation operations. Specifically, one is to adopt the addition/concatenation operation, PA and CA in serial to fuse the outputs of the local patch-based denoising process and GCL module, and the other is to adopt the addition/concatenation operation before PA and the concatenation/addition operation before CA, which are in parallel. As exhibited in Table IV, the feature fusion module can better fuse these two outputs by using our fusion mode.
TABLE IV: Ablation study of different fusion modes in the feature fusion module.
Fusion Mode PSNR SSIM CIEDE2000
addition PA CA in serial 26.40 0.895 4.312
concatenation PA CA in serial 27.49 0.900 4.243
concatenation PA, addition CA in parallel 27.68 0.899 4.329
addition PA, concatenation CA in parallel (ours) 27.91 0.900 3.945
Effect of sampling steps in sampling process. During the sampling process, diffusion model iteratively denoises until generating sharp results through $S$ steps. Theoretically, a large sampling step contributes to producing more visually pleasing result [29], whereas it also increases the processing time. To further explore the performance of different sampling steps, we perform an ablation experiment setting different $S$ from 5 to 25. Table V exhibits the average PSNR, SSIM and CIEDE2000 of different sampling steps. As reported, we can find that the performance degrades slightly as the number of sampling step is greater than 10. To carefully balance the performance and sampling time, we set $S$ to 10 in our experiment.
TABLE V: Ablation study on the sampling steps in sampling process (hyperparameter $S$)
@@
Effect of the mask scale in FIR. The mask scale (hyperparameter $\beta$) is the weight of different amplitude components in FIR. Obviously, when the hyperparameter $\beta$ is set to 0, the Fourier-aware iterative refinement module is invalid. When $\beta$ is set to 1, FIR focuses on fusing the phase component of forward process and amplitude component of reverse process in Fourier domain. Table VI shows the dehazing performance of different scale values from 0.01 to 1. As reported, FIR can better help diffusion model to restore haze-free images when the $\beta$ is set to 0.1.
TABLE VI: Ablation study on the mask scale in FIR (hyperparameter $\beta$)

### v2

@@ IV-C Global Compensated Learning
It is significant that the image restoration methods can handle unlimited-size images in practice. To achieve unlimited-size image restoration, we adopt the patch-based processing strategy that divides images into multiple fixed-size patches and then deals with each patch independently. However, this patch-based strategy may produce artifacts with the ignorance of global dependency information in all patches. To address this, we propose an effective GCL module that adopts FFB as the basic U-net architecture unit to capture all patches’ global dependency information. After learning global feature representations, the designed fusion module fuses the local patch-based image restoration process and GCL module in a pixel-channel attention collaborative manner. The structure of GCL module and the fusion module are illustrated in Fig. 2.
@@ IV-D Phased Training Strategy
Remark: The noise-estimation constraint is necessary for the local conditional diffusion model. In the reverse process, the noise subtracted at each time step is estimated by the U-Net network. The network can predict the noise $\epsilon_{0}\left(x_{t},\bar{x},t\right)$ with the noise-estimation constraint (Eq. 5), further obtaining the mean and variance of $x_{t-1}$ (Eq. 4). If we remove the noise-estimation constraint in the early stages (i.e., only the reconstruction constraints used in the training stages of local conditional diffusion model), the noise prediction network is invalid, causing severe degradation of sampling quality. In this case, only using the reconstruction constraints makes it difficult to predict the dehazed images from input images with Gaussian noise distribution. When we only use the reconstruction constraints in the whole training stages, the dehazed results are pure noisy images even training for 30w iterations in the ERICE dataset. It further validates our analysis.
Apart from the above mentioned loss function that is optimized over local patch, we also exploit the perceptual loss to global compensated learning module for constraining the whole image. The perceptual loss function can measure high-level perceptual similarity between the output and ground truth by extracting the multi-layer features from pre-trained network VGG-19 [57]. By minimizing perceptual loss, the restored output $f_{2}$ has more similar texture and structure features with ground-truth images. Similar to [58], we hierarchically extract five latent feature maps from VGG-19 model to measure the image perceptual similarity between the output $f_{2}$ and ground-truth images $z$, which can be expressed as:
@@
where $F_{i}\left(\cdot\right)$ represents the $i$-th latent features obtained by VGG-19. $\omega_{i}$ denotes the weight coefficient of the $i$-th latent features. Here, we obtain latent features from layers 1, 6, 11, 20, and 29 in the VGG-19 model, and the $\omega_{i}$ ($i$ from 1 to 5) are set to $\frac{1}{32},\frac{1}{16},\frac{1}{8},\frac{1}{4},1$, respectively. Furthermore, we also employ pixel-level $\mathcal{L}_{1}$ loss to assess the diversity between $f_{2}$ and $z$. Hence, the loss function $\mathcal{L}_{global}$ in global compensated learning branch is shown as:
@@ V Experiments
In this section, the implementation details of our experiments are first described. Next, the comparisons with several recent approaches on the synthetic and real-world datasets are performed. After that, a spectral consistency analysis is conducted to verify the spectral similarity between restored images and reference images. Then, we conduct ablation studies to demonstrate the effectiveness of core components in RSHazeDiff, followed by the application of our method in intelligent transportation systems. Finally, we discuss the limitation of our method and future work.
@@ V-A Implementation Details
Dataset. We evaluate the performance of RSHazeDiff on three RS hazy datasets, i.e., the light hazy image dataset (LHID), the dense hazy image dataset (DHID) [21] and expanded remote sensing image cloud removing dataset (ERICE) [21]. LHID and DHID are synthetic datasets, and each image has 512$\times$512 pixels. LHID contains 30,517 image pairs for training and 1000 for testing (LHID-A and LHID-B). The DHID dataset contains 14,490 image pairs for training and 500 for testing. The DHID dataset is generated by randomly sampling the ambient light $A\in\left[0.7,1\right]$ and 500 different real-world transmission maps. The ERICE is a real-world dataset expanded on RICE1 [59], collected in some scene categories from Google Earth, such as mountain, river, and desert. In ERICE dataset, 1,850 and 150 images are regarded as training and test sets, each with a resolution of 256$\times$256 pixels. Additionally, we employ the NHM dataset [60] to evaluate the nighttime dehazing capability of RSHazeDiff. Like [60], the training set is 8,073 image pairs from NHR dataset[60].
TABLE I: Quantitative comparisons with 17 dehazing approaches on the synthetic LHID and DHID datasets. The $\uparrow$ means higher is better while $\downarrow$ means lower is better. The red and blue colors are used to indicate the $1^{st}$ and $2^{nd}$ ranks, respectively
@@
DCP [8] TPAMI’10 21.12 0.818 9.161 0.481 21.49 18.92 0.824 9.867 0.163 79.92
AOD-Net [14] ICCV’17 22.24 0.842 7.122 0.475 32.61 15.37 0.718 14.272 0.269 151.16
GridDehazeNet [17] ICCV’19 26.64 0.883 4.597 0.451 18.86 25.51 0.868 5.053 0.121 65.95
IDeRs [10] INS’19 14.68 0.627 19.563 0.633 44.65 13.33 0.575 19.942 0.344 120.17
FFA-Net [54] AAAI’20 27.79 0.887 4.424 0.451 16.27 25.05 0.880 5.590 0.144 73.27
PhysicsGAN [25] TPAMI’20 22.01 0.835 8.198 0.512 38.72 19.85 0.824 10.415 0.187 94.61
PSD [18] CVPR’21 23.33 0.855 6.998 0.549 37.25 23.30 0.853 6.911 0.163 82.66
FCTF-Net [20] GRSL’21 25.85 0.876 5.423 0.459 21.21 17.21 0.739 12.794 0.221 128.49
DCIL [21] TGRS’22 28.12 0.898 5.180 0.447 14.93 26.95 0.896 4.877 0.131 64.73
SGID-PFF [22] TIP’22 24.12 0.852 7.091 0.488 30.13 25.69 0.867 6.147 0.158 80.74
Trinity-Net [34] TGRS’23 24.78 0.867 6.737 0.460 17.76 25.86 0.878 5.479 0.139 75.46
RIDCP [61] CVPR’23 26.75 0.843 4.553 0.128 19.07 26.15 0.874 5.329 0.136 64.36
C2PNet [62] CVPR’23 28.65 0.867 4.065 0.086 15.16 26.20 0.871 4.925 0.106 56.40
Focal-Net [19] ICCV’23 29.39 0.899 4.012 0.445 14.32 25.73 0.887 5.313 0.131 73.04
PSMB-Net [63] TGRS’23 28.89 0.900 5.864 0.454 16.13 27.07 0.887 4.965 0.113 53.83
AU-Net [64] TGRS’24 28.44 0.893 4.045 0.451 16.96 26.42 0.891 4.489 0.138 69.28
Ours - 29.65 0.905 3.875 0.436 12.85 27.91 0.900 3.945 0.102 49.52
@@
Training Details. All experiments are implemented by PyTorch [65] framework with NVIDIA GeForce RTX 3090 GPU. In the training of Fourier-aware conditional diffusion model, we adopt the Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$) optimizer with a batch size of 4. The initial learning rate is set to 0.00002. The single input image is randomly cropped to 16 patches for training, each patch with a resolution of 64$\times$64. We adopt PTS to train the model until it converges. Concretely, we first train the model with noise estimation constraint for 1,500,000 iterations, and then refine it with reconstruction constraint for 500,000 iterations in LHID. Similarly, we first train the model for 1,000,000 iterations and then refine it for 500,000 iterations in DHID. To expedite the sampling process, we adopt the implicit sampling strategy [66] in which the total number of time steps and implicit sampling steps are set to 1,000 and 10. Furthermore, we use the exponential moving average (EMA) strategy [67] to eliminate slight color shift problem. During inference, the single input image is decomposed into $D$ overlapping patches, each with a resolution of 64$\times$64. $D$ denotes the number of overlapping patches and is set to 64. For the training of the global compensated learning module, we use the Adam optimizer with the batch size of 8 and the initial learning rate of 0.001.
@@
Evaluation Settings. To demonstrate the superiority of RSHazeDiff, we qualitatively and quantitatively compare it against 16 representative dehazing algorithms, including DCP [8], AOD-Net [14], GridDehazeNet [17], IDeRs [10], FFA-Net [54], PhysicsGAN [25], PSD [18], FCTF-Net [20], DCIL [21], SGID-PFF [22], Trinity-Net [34], RIDCP [61], C2PNet [62], Focal-Net [19], PSMB-Net [63], and AU-Net [64].
@@
We adopt PSNR [68], SSIM [68], and CIEDE2000 [69] as the distortion metrics and LPIPS [70] and FID [71] as the perceptual metrics for dehazing performance evaluation. PSNR index assesses the error between the dehazed images and the corresponding ground-truth ones pixel by pixel. SSIM measures image similarity regarding brightness, contrast, and structure. CIEDE2000 indicator is employed to evaluate the color distinction between the dehazed image and its haze-free counterpart. LPIPS and FID assess the perceptual quality and realism of the dehazed images. Larger values of PSNR and SSIM indicate better results, while smaller values of CIEDE2000, LPIPS, and FID indicate better performance.
@@ V-B Comparison with State-of-the-art Methods
Results on Synthetic Dataset. Table I reports the quantitative performance of 17 dehazing algorithms on synthetic LHID and DHID test datasets. To make fair comparisons, we re-train all compared methods on synthetic datasets according to the implementation details of their papers. Note that we directly re-train the whole network of DCIL on DHID dataset without the phased learning strategy [21]. As observed, our RSHazeDiff can achieve the highest PSNR and SSIM values and the lowest CIEDE2000 and FID values than other representative dehazing algorithms, which indicates that our RSHazeDiff is able to better restore the texture details and color information from hazy RS images. Moreover, Table II offers comparisons with several nighttime dehazing methods on the NHM dataset. As observed, RSHazeDiff wins first and second place among ten representative nighttime dehazing methods, revealing its potential in nighttime dehazing tasks.
TABLE II: Quantitative comparisons on the nighttime hazy dataset NHM
[…]
NDIM [72] ICIP’14 12.69 0.617
GS [73] ICCV’15 11.85 0.615
MRP [74] CVPR’17 13.11 0.655
FAST-MRP [74] CVPR’17 13.31 0.649
OSFD [60] MM’20 13.28 0.667
VD [75] CVPRW’22 13.77 0.687
GHLP [76] TIP’22 13.15 0.674
GAPSF [77] MM’23 13.48 0.665
NightHazeFormer [36] MM’23 18.54 0.784
Ours - 15.86 0.786
@@
Fig. 5 presents the visual comparisons against several recent dehazing approaches on DHID datasets. We can observe that AOD-Net tends to darken the whole images and produce noticeable Halo artifacts, degrading the overall visibility of the dehazed images. FFA-Net and Focal-Net cannot remove haze effectively when dealing with dense hazy images, leading to some haze residuals. In addition, the dehazing results of FFA-Net suffer from slight color distortion. Although DCIL can remove a majority of haze to a certain extent, it may distort some texture details of input images and fails to completely remove the haze in some regions. Trinity-Net cannot produce pleasing dehazing results due to haze residuals and serious color distortion. Compared with above dehazing algorithms, our RSHazeDiff can generate much clearer haze-free images with realistic color fidelity and finer texture details.
@@
Results on Real-World Hazy Images. To further validate the dehazing performance on real-world images, we evaluate our method against several recent dehazing approaches on ERICE dataset. The ERICE dataset contains some images captured in thick haze scenarios. Table III exhibits the quantitative comparison results of different dehazing algorithms. As reported, the proposed method surpasses other dehazing algorithms in terms of PSNR and SSIM, indicating that the images restored by RSHazeDiff are most similar to its corresponding ground-truth images. Although CIEDE2000 value of RSHazeDiff is not the best, our method wins the second place among eleven competitive dehazing algorithms. The quantitative results verify that the superiority of RSHazeDiff on real-world RS images.
Remark: The above experiment applies the model trained on the training set of ERICE dataset to evaluate the performance of real-world haze removal. To further evaluate the generalization ability of our model, we train the model on the synthetic dataset DHID and test it on an unseen real-world dataset ERICE. Compared to Table III, all methods suffer from obvious performance degradation, typically because of the domain gap between synthetic and real-world hazy images. Comfortingly, our method wins first place in these indicators. To avoid repetition, we just exhibit the results in Table III.
@@
TABLE III: Average PSNR, SSIM and CIEDE2000 of various different approaches on the real-world ERICE dataset
@@
Generalization to out-of-distribution degradations. To show the generalization ability of RSHazeDiff, we train our model on the synthetic dataset DHID and test it on the training-unseen datasets D-Hazy [78] and T-Cloud [79]. The two datasets have different haze distribution and climatic conditions from the training set. The quantitative evaluation results are shown in Table IV. As reported, we win first place in terms of PSNR, SSIM, and CIEDE2000, and also achieve impressive performance in terms of LPIPS and FID, which further verifies the favorable generalization ability of our method.
TABLE IV: Quantitative comparisons with recent dehazing approaches on the D-Hazy and T-Cloud test set
Method D-Hazy T-Cloud
[…]
AOD-Net [14] 13.55 0.648 15.060 0.262 41.39 19.32 0.602 11.974 0.317 111.38
FFA-Net [54] 11.48 0.648 19.817 0.314 34.20 19.27 0.619 12.637 0.392 138.79
DCIL [21] 11.03 0.686 19.801 0.304 33.23 20.08 0.716 10.590 0.369 145.43
RIDCP [61] 12.20 0.677 18.778 0.311 34.70 19.71 0.661 10.989 0.338 114.39
Focal-Net [19] 8.92 0.608 26.116 0.409 40.59 18.57 0.655 11.364 0.352 116.60
PSMB-Net [63] 13.14 0.636 18.641 0.447 79.28 20.05 0.709 16.861 0.457 122.37
AU-Net [64] 13.12 0.717 15.351 0.289 37.30 19.52 0.692 11.903 0.400 135.40
Ours 13.66 0.725 15.008 0.236 39.75 20.35 0.739 10.444 0.337 114.54
@@ V-D Ablation Study
Effect of different components in RSHazeDiff. Our method shows dehazing superiority both quantitatively and qualitatively over 17 representative dehazing methods. Furthermore, we perform an ablation study to analyze the effectiveness of the proposed components, including the phased training strategy, Fourier-aware iterative refinement, and global compensated learning module.
@@
For fair comparisons, all variants are trained in the same way as before and tested on DHID dataset. The performances of these variants can be found in Table V and Fig. 1.
TABLE V: Ablation study of different components on DHID dataset.
@@
As shown in Table V and Fig. 1, each component of RSHazeDiff can boost the dehazing performance. When applying the phased training strategy, the model achieves advanced performance that has fewer haze residuals and higher metric values over baseline model. Furthermore, the introduction of FIR and GCL module also have brought performance improvement of model in three indicators. In short, the full model performs favorably against the other variants both quantitatively and qualitatively, demonstrating that all the components of RSHazeDiff are beneficial and necessary for effective haze removal.
Effect of sampling steps in sampling process. During the sampling process, diffusion model iteratively denoises until generating sharp results through $S$ steps. Theoretically, a large sampling step contributes to producing more visually pleasing result [30], whereas it also increases the inference time. To further explore the performance of different sampling steps, we perform an ablation study setting different $S$ from 5 to 25. Table VI exhibits the average PSNR, SSIM and CIEDE2000 of different sampling steps. As reported, we can find that the performance degrades slightly as the number of sampling step is greater than 10. To carefully balance the performance and sampling time, we set $S$ to 10 in our experiment.
TABLE VI: Ablation study on the sampling steps in sampling process (hyperparameter $S$)
@@
Effect of the mask scale in FIR. The mask scale (hyperparameter $\beta$) is the weight of different amplitude components in FIR. Obviously, when the hyperparameter $\beta$ is set to 0, the Fourier-aware iterative refinement module is invalid. When $\beta$ is set to 1, FIR focuses on fusing the phase component of forward process and amplitude component of reverse process in Fourier domain. Table VII shows the dehazing performance of different scale values from 0.01 to 1. As reported, FIR can better help diffusion model to restore haze-free images when the $\beta$ is set to 0.1.
TABLE VII: Ablation study on the mask scale in FIR (hyperparameter $\beta$)
@@
[…]
TABLE VIII: Performance comparisons of different global modeling-based methods on the ERICE dataset
TransWeather [81] Fourmer [82] FSNet [83] Ours
PSNR 33.92 32.12 34.67 36.56
SSIM 0.934 0.932 0.941 0.958
LPIPS 0.053 0.074 0.059 0.052
FID 41.66 59.29 36.94 32.97
V-E Application
RS image dehazing provides clear observation of transportation infrastructure, such as roads, vehicles, tunnels, bridges, and buildings. That enables downstream visual-based transportation tasks to extract useful traffic information from images under hazy conditions. The downstream transportation applications include road extraction, urban road planning, traffic congestion monitoring, vehicle detection, infrastructure development, and disaster response. Here we take vehicle detection as an example to exhibit the applicability of our approach in transportation field.
Vehicle detection helps road network planning and transport traffic statistics, widely used in vision-based intelligent transportation systems, yet its detection accuracy is noticeably degraded under hazy conditions. To verify that our RSHazeDiff can benefit vehicle detection, we employ a pre-trained YOLOv5s [84] detector to detect the vehicles on the original hazy images and the dehazing results by different methods. The YOLOv5s detector is retrained on the fine-grained vehicle detection dataset VEDAI [85] to ensure stable vehicle detection. Fig. 10 exhibits two samples from the dense hazy images DHID. As shown, after dehazing, the detection performance is significantly improved. Compared to other dehazing algorithms, our method succeeds in removing thick haze with sharper structures and details, and achieves superior accuracy for vehicle detection.
Figure 10: Vehicle detection results on the hazy images and images after dehazing by different methods. From (a) to (e): (a) vehicle detection results on the hazy images, and the vehicle detection results after dehazing by (b) C2PNet [62], (c) Focal-Net [19], (d) PSMB-Net [63], and (e) our RSHazeDiff, respectively.
V-F Limitation and Future Work
Although RSHazeDiff achieved impressive performance, our method requires longer inference time compared to end-to-end image dehazing networks, due to the multiple iterative evaluations and gradient calculations in DDPM. Even if we adopt the DDIM-based sampling strategy to reduce the number of sampling steps to 10, the inference speed of our method is still slower than the end-to-end dehazing approaches. Our RSHazeDiff takes 1.016 seconds to restore an image of size 256 × 256 on a single NVIDIA GeForce RTX 3090 GPU. Recently, the latent diffusion model [39] performs the diffusion process in a compressed latent space rather than the pixel space. It will be a potential direction to improve the computational efficiency of our method. In future work, we will explore the dehazing performance of diffusion model in the lower dimension latent space.